# An Autonomous Vision-Based Algorithm for Interplanetary Navigation

### v2

@@ 4.2 Measurement Model
One of the contributions of the work is to present a novel measurement model for deep-space triangulation. The exploited measurement model expresses the position of the observations in pixel coordinates in the camera reference frame $\mathbb{C}$. Moreover, it embeds the light effects and their dependencies, which are described with respect to the planet and spacecraft state.
In the existing literature, external observations are frequently modeled as LoS directions or angular coordinates, such as azimuth and elevation [8, 19, 42]. This conventional approach involves manipulating raw camera measurements (i.e., pixel coordinates), resulting in a non-Gaussian error distribution compared to the original raw data. Moreover, prior studies have attempted non-linear correction of light-aberration effects on the observations by employing the a priori estimate of velocity. However, this method couples measurement and process noise, thus violating the Kalman filter assumption [19, 9]. Furthermore, light-time delay is typically evaluated in the literature using an iterative approach, which increases the computational time required to achieve an optimal solution.
In contrast, the developed measurement model utilizes pixel coordinates to represent observations and incorporates light effects within the model itself to enhance correlations and improve linearity, thus avoiding coupling between measurement and process noises. Eventually, the light-time delay is derived through a linear approximation of the equation describing the light-time effect.
@@ 4.2.1 Evaluation of the time delay
Eq. (32) is a constraint that links the spacecraft state with the planet position. Moreover, as the planet motion does not have an analytical solution, the value of $\tau$ cannot be solved analytically to implicitly include this effect in the measurement. Yet, it is possible to linearize the constraint in Eq. (32) with respect to $\Delta t$ under the assumption that the time delay $\Delta t=t-\tau$ is small with respect to the orbital period. Thus, the planet motion can be approximated by a linear motion with constant velocity as follows:
@@
$\mathfrak{L}$ $\simeq c^{2}\Delta t^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\left(\bm{r}_{\rm{pl}}(t)-\bm{v}_{\rm{pl}}(t)\Delta t-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t)-\bm{v}_{\rm{pl}}(t)\Delta t-\bm{r}(t)\right)$ (34)
@@
Eq. (35) shows that two solutions are possible, given the geometry between the planet and the spacecraft. It is important to understand which solution is the correct one to uniquely solve for $\Delta t$. By defining $\bm{\beta}_{\rm{pl}}(t)=\dfrac{\bm{v}_{\rm{pl}}(t)}{c}$ and the angle between $\bm{r}_{\textrm{pl/sc}}(t)$ and $\bm{v}_{\textrm{pl}}(t)$ as the planet flight path angle $\epsilon$, the approximated solution for the light-time correction is:
@@
$\Delta t=\frac{\left|\left|\bm{r}_{\rm{pl/sc}}(t)\right|\right|}{c\left(1-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\right)}\left(-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|\cos\epsilon+\sqrt{\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{2}}\left(\cos^{2}\epsilon-1\right)+1}\right)$ (37)
Since $\left|\left|\bm{\beta}_{\textrm{pl}}(t)\right|\right|\leq 1$, $1-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\geq 0$ $\forall\bm{\beta}_{\rm{pl}}(t)$ and $\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\left(\cos^{2}\epsilon-1\right)+1\geq 0$ $\forall\bm{\beta}_{\rm{pl}}(t)$ and $\forall\epsilon$. Note that $\cos\epsilon\geq 0$ $\forall\epsilon$ by flight path angle definition. Eq. (37) provides an analytical solution at first order for the light-time delay which can be exploited to include light-time correction in the filter update.
@@ 4.2.2 Definition of the measurement model equation
$\bm{l}_{\rm{pl/sc}}=\frac{\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|}$ (38)
@@
${}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\frac{1}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(3)}}\begin{pmatrix}{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(1)}\\ {}_{h}^{\mathbb{C}}{r}_{\rm{pl},(2)}\end{pmatrix}$ (41)
where ${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ is the projection of the planet line of sight in the image plane in homogeneous coordinates, ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ represents the planet location in pixel coordinates, ${}_{h}^{\mathbb{C}}{r}_{\rm{pl},(i)}$ is the $i$-esimal coordinate of vector ${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}$, $\bm{K}_{\rm{cam}}$ is the camera intrinsic matrix, and $\bm{A}_{\rm{corr}}$ is the rotation matrix corrected for the stars light aberration from the inertial reference frame $\mathcal{I}$ to the camera reference frame $\mathcal{C}$.
@@ 4.2.3 Definition of the Jacobian of the measurement model
In this section the measurement model derivatives with respect to the filter state and all other possible sources of errors are computed by variational analysis. The considered error sources include the spacecraft attitude and planet state. Once variations are computed, it is possible to use the derivatives with respect to the state to construct the measurement model Jacobian matrix. Moreover, by using the derivatives of the measurement with respect to the other error sources, it is possible to inflate the measurement covariance matrix to account for uncertainties in the spacecraft attitude and planet ephemerides.
To perform the variational analysis, the variation of each parameter is considered, and its nonlinear mapping is approximated with a linear function leveraging its first derivatives. This leads to determining the mapping between the variation of the output parameter with respect to the variation of the parameters under analysis. Note that $\delta(\cdot)$ is the variation of the considered parameter. Thus:
@@
$\delta{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K}_{\rm{cam}}\bm{A}_{\rm{corr}}\left(\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}+2\left[{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}\right]^{\wedge}\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}\right)$ (43)
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}$ is the vectorial part of the quaternion $\bm{q}_{\mathcal{C}/\mathcal{I}}$ representing the rotation from the inertial reference frame $\mathcal{I}$ to the camera reference frame $\mathcal{C}$.
@@
$\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}=$ $\left(\bm{I}_{3x3}+2\,{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}^{\top}\right)\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}+$ (44)
$+\left({}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{3x3}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}\right)\delta{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}$
@@
Finally, by combining Eqs. (42)–(47), the linear mapping between the variation of ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ and the variation of $\bm{r}_{\rm{pl}}$, $\bm{v}_{\rm{pl}}$, $\bm{r}$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}$ can be established:
$\delta{}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ $=\begin{bmatrix}\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{r}(t)}&\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{v}(t)}&\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}}&\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{r}_{\rm{pl}}(t)}&\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{v}_{\rm{pl}}(t))}\end{bmatrix}\begin{pmatrix}\delta\bm{r}(t)\\ \delta\bm{v}(t)\\ \delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}\\ \delta\bm{r}_{\rm{pl}}(t)\\ \delta\bm{v}_{\rm{pl}}(t)\end{pmatrix}=$ (48)
$=\begin{bmatrix}\dfrac{1}{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(3)}}&0&-\dfrac{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(1)}}{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(3)}^{2}}\\ 0&\dfrac{1}{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(3)}}&-\dfrac{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(2)}}{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(3)}^{2}}\end{bmatrix}\bm{K}_{\rm{cam}}\bm{A}_{\rm{corr}}\begin{bmatrix}\bm{R}_{\rm sc}&\bm{V}_{\rm sc}&\bm{Q}_{\rm att}&\bm{R}_{\rm pl}&\bm{V}_{\rm pl}\end{bmatrix}\begin{pmatrix}\delta\bm{r}(t)\\ \delta\bm{v}(t)\\ \delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}\\ \delta\bm{r}_{\rm{pl}}(t)\\ \delta\bm{v}_{\rm{pl}}(t)\end{pmatrix}$
where matrices $\bm{R}_{\rm sc}$, $\bm{V}_{\rm sc}$, $\bm{Q}_{\rm att}$, $\bm{R}_{\rm pl}$, and $\bm{V}_{\rm pl}$ are reported in Appendix A.
Eq. (48) provides the linear mapping between the projection of the planet in the image plane as a function of the spacecraft state, the spacecraft inertial attitude, and planet ephemeris. Note that matrices $\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{r}(t)}$ and $\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{v}(t)}$ are used to build the measurement model Jacobian $\bm{H}$ employed in the Extended Kalman Filter. Moreover, $\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}}$, $\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{r}_{\rm{pl}}(t)}$, and $\frac{\partial{}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{v}_{\rm{pl}}(t))}$ may be used to inflate the measurement covariance matrix taking into account attitude determination uncertainty from the attitude filter and the planet ephemeris uncertainties induced by an onboard implementation of the planet motion.
Note that the advantages of the proposed measurement model are threefold. First, the centroid measurement from the camera raw image is directly provided to the filter without any calculation. Indeed, any operation could increase the non-linearity of the measurement error distribution, leading to less linear behavior of the filter posterior estimates. This is particularly true when trigonometric functions are used as in previous work [9, 19]. Second, the expression found in Eq. (37) is analytic, leading that no optimization is necessary on board to compute the light time delay. This is in contrast with state-of-the-art techniques that use Newton-Raphson’s method to find the light-time correction [45, 9, 19]. Third, as the light time correction is embedded in the measurement model, uncertainties arising from the planets ephemeris and correlations with spacecraft state can be taken into account as shown in Eq. (48).
When only light-time correction is included in the measurement model (as in Section 5 for comparison analysis), Eqs. (39) and (44) are not used and Eqs. (40) and (43) simply become:
${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K_{\textrm{cam}}}\bm{A}_{\rm{corr}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}$ (49)
$\delta{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K_{\textrm{cam}}}\bm{A}_{\rm{corr}}\left(\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}+2\left[{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}\right]^{\wedge}\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{q}_{v}}\right)$ (50)

### v3

@@ 4.2 Measurement Model
One of the contributions of the work is to present a novel measurement model for deep-space triangulation. In the state of the art the LoS measurements have been modeled as azimuth and elevation observations [9, 19, 38]. Thus, the raw camera measurement is manipulated to obtain a derived quantity to be included in the filter. This implies that the error distribution is less Gaussian with respect to the raw measurement error distribution. Moreover, the correction of the light-aberration effect is performed on the observation through the exploitation of the estimated velocity information, coupling the measurement and the process noise. On the contrary, the measurement model proposed hereafter expresses the observations in pixel coordinates in the camera plane. In addition, it embeds the light effects and their dependencies with respect to the planet and spacecraft state. Therefore, the navigation filter takes into account these effects during the mean and covariance update, and the coupling between the measurement and the process noise is avoided.
@@ 4.2.1 Evaluation of the time delay
Eq. (26) is a constraint that links the spacecraft state with the planet position. Moreover, as the planet motion does not have an analytical solution, the value of $\tau$ cannot be solved analytically to implicitly include this effect in the measurement. Yet, it is possible to linearize the constraint in Eq. (26) with respect to $\Delta t$ under the assumption that the time delay $\Delta t=t-\tau$ is small. The planet motion can be approximated linearly as:
@@
$\mathfrak{L}$ $\simeq c^{2}\Delta t^{2}+\left(\bm{r}_{\rm{pl}}(t)-\bm{v}_{\rm{pl}}(t)\Delta t-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t)-\bm{v}_{\rm{pl}}(t)\Delta t-\bm{r}(t)\right)=$ (28)
@@
Eq. (29) shows that two solutions are possible, given the geometry between the planet and the spacecraft. It is important to understand which solution is the correct one to uniquely solve for $\Delta t$. By defining $\bm{\beta}_{\rm{pl}}(t)=\dfrac{\bm{v}_{\rm{pl}}(t)}{c}$ and the planet flight path angle $\epsilon$, the approximated solution for the light-time correction is:
@@
$\Delta t=\frac{\left|\left|\bm{r}_{\rm{pl/sc}}(t)\right|\right|}{c\left(1-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\right)}\left(-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|\cos\epsilon+\sqrt{\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|\left(\cos^{2}\epsilon-1\right)+1}\right)$ (31)
It is worthy noting $1-\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\geq 0$ $\forall\bm{\beta}_{\rm{pl}}(t)$ and $\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|^{2}\left(\cos^{2}\epsilon-1\right)+1\geq 0$ $\forall\bm{\beta}_{\rm{pl}}(t)$ and $\forall\epsilon$, as $\left|\left|\bm{\beta}_{\rm{pl}}(t)\right|\right|\leq 1$. Note that $\cos\epsilon\geq 0$ $\forall\epsilon$ by flight path angle definition. Eq. (31) provides an analytical solution at first order for the light-time delay which can be exploited to include light-time correction in the filter update.
@@ 4.2.2 Definition of the measurement model equation
$\bm{l}_{\rm{pl/sc}}=\frac{\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)}{\left|\left|\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)^{\top}\left(\bm{r}_{\rm{pl}}(t-\Delta t)-\bm{r}(t)\right)\right|\right|}$ (32)
@@
${}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\frac{1}{{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(3)}}\begin{pmatrix}{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(1)}\\ {}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(2)}\end{pmatrix}$ (35)
where ${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ is the projection of the planet line of sight in the image plane in homogeneous coordinates, ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ is the same vector but in non-homogeneous coordinates, ${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl},(i)}$ is the $i$-esimal coordinate of vector ${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}$, $\bm{K}_{\rm{cam}}$ is the camera intrinsic matrix, and $\bm{A}_{\rm{corr}}$ is the rotation matrix from the inertial reference frame $\mathcal{N}$ to the camera reference frame $\mathcal{C}$ corrected for the stars light aberration.
@@ 4.2.3 Definition of the Jacobian of the measurement model
As the measurement model is analytic, its Jacobian can be easily computed by variational analysis. Thus:
@@
$\delta{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K}_{\rm{cam}}\bm{A}_{\rm{corr}}\left(\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}+2\left[{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}\right]^{\wedge}\delta\bm{q}_{\mathcal{C}/\mathcal{N}}^{(v)}\right)$ (37)
where $\bm{q}_{\mathcal{C}/\mathcal{N}}^{(v)}$ is the vectorial part of the quaternion representing the rotation from the inertial reference frame $\mathcal{N}$ to the camera reference frame $\mathcal{C}$.
@@
$\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\rm{aberr}}=$ $\left(\bm{I}_{3x3}+2\,{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}^{\top}\right)\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}+$ (38)
$+\left({}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}-{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}^{\top}\right)\delta{}^{\mathcal{}}\bm{\beta}_{\rm{sc}}$
@@
Finally, by combining Eqs. (36)–(41), the linear mapping between the variation of ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ and the variation of $\bm{r}_{\rm{pl}}$, $\bm{v}_{\rm{pl}}$, $\bm{r}$, and $\bm{q}_{\mathcal{C}/\mathcal{N}}^{(v)}$ can be established:
$\delta{}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{\Pi}\begin{pmatrix}\delta\bm{r}(t)\\ \delta\bm{v}(t)\\ \delta\bm{q}_{\mathcal{C}/\mathcal{N}}^{(v)}\\ \delta\bm{r}_{\rm{pl}}(t)\\ \delta\bm{v}_{\rm{pl}}(t)\end{pmatrix}$ (42)
Eq. (42) provides the linear mapping between the projection of the planet in the image plane as a function of the spacecraft state and of the planet motion in the solar system. In an operational scenario, the variation of the spacecraft state comes from the orbital and attitude filter performance, whereas the variation of the planet motion arises from the knowledge of the planet ephemeris or from the approximation errors induced by an onboard implementation of the planet motion. The Jacobian matrix of the measurement model exploited in the navigation filter takes into account the first six columns of $\bm{\Pi}$, being only the probe position and velocity contained in the state vector $\bm{x}$.
The advantages of the proposed measurement model are threefold. Firstly, the proposed model avoids any operation on the raw measurement extracted from the deep-space images to correct the light-aberration effect. Secondly, the expression found in Eq. (31) is analytic and provides a first-order approximation for the light-time delay which can be used to find efficiently on board $\Delta t$ without any optimization, at the contrary of [9]. Thirdly, the proposed equation models the light effects as a function of the spacecraft and planet states, implying that their uncertainty can be taken into account during filter updates.
In the following, two approaches to correct the light-aberration effect are analyzed. Case 1 exploits the complete measurement model developed in this section. Whereas, in case 2 only the light-time correction is included in the measurement model, as the relativistic light aberration is compensated when performing attitude determination on all the bright objects in the image (see Sec. 3). This implies that Eqs. (33) and (38) are not used and that Eqs. (34) and (37) simply become:
${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K}\bm{A}_{\rm{corr}}{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}$ (43)
$\delta{}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\bm{K}\bm{A}_{\rm{corr}}\left(\delta{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}+2\left[{}^{\mathcal{}}\bm{l}_{\rm{pl/sc}}\right]^{\wedge}\delta\bm{q}_{\mathcal{C}/\mathcal{N}}^{(v)}\right)$ (44)